Date
Name
Other Observers
Instrument
Hours Observed
Hours Lost
Reason
Observing Program
12/04/2021
Nuria Torres-Albà
Scott, Dieter
SARA CT
7
0
No problems to report
Clemson University observations of high-z blazars No problems to report
12/03/2021
William Keel
AY203 coass
Andor CCD #2
8
0
No problems to report
Galaxies and H II regions. Focus 127200, guider focus 0.
11/30/2021
Tom Rutherford
Gary Henson
CCD
6
0
No problems to report
Asteroid light curve.
11/28/2021
Tom Rutherford
None
CCD
7
0
No problems to report
Exoplanet transit.
11/27/2021
Todd Hillwig
None.
Andor CCD
8
0
No problems to report
PN central star monitoring.
11/24/2021
Tom Rutherford
None
CCD
6
0
No problems to report
Standard Stars.
11/23/2021
Tom Rutherford
None
CCD
8
0
No problems to report
Exoplanet. No issues.
11/23/2021
Priya Gokuldass
Kamp
ccd
2
8
Clouds
TESS follow up observations. No new issues to report.
11/22/2021
Paulina Karczmarek
None
Andor CCD
9
0
No problems to report
Observing program: absolute photometry in Sloan bands for local comparison stars of bright variables (Cepheids, RR Lyrae).
11/19/2021
Ross Silver
Dieter, Yong, Isaiah
CCD
8
0
No problems to report
Obtaining photometric redshift of blazars.
11/18/2021
Paulina Karczmarek
None
Andor CCD
8
0
Equipment Failure
Observing program: absolute photometry in Sloan bands for local comparison stars of bright variables (Cepheids, RR Lyrae). Dropout shutter would not open entirely and indicated "opening" all the time. I don't know if this is a software problem or actual problem with the shutter engine.
11/18/2021
Tom Rutherford
None
CCD
7
0.5
Software/Control PC Issues
Asteroid Photometry. Lower shutter gave no indication that it was open so I kept it closed (which meant that I could not reach one of my low targets). Could not access the telescope computer when I first logged in, but cycling the APC power rack brought it back up. All worked fine after that.
11/18/2021
Albert, Icarus
none
n/a
0
6
Connection Issues
Could not connect to telescope camera
11/18/2021
Albert, Icarus
Emmett Dalton
Bessel IRVB filters
6
0
No problems to report
data collection
11/14/2021
Gillam
Lilyanne Pepe, Aakash Thiyagarajan
Andor Camera
11
0
Equipment Failure
telescope control system indicated the dropout door hadn't opened all the way. We operated with that door closed. rOAp stars
11/12/2021
Keefe Kamp
Van Putten, Corcoan
CCD
7
1
Equipment Failure
Dec and RA limits triggered at login (Cardinal directions displayed in limit boxes). Had to have Peter fix it. Dropout would not detect open however would be open. Note on notepad discusses how to open. Other than that no other problem occurred. TESS Follow up observations done.
11/11/2021
Emmett Dalton
CCD
1
4
Equipment Failure
Dropout shutter still not opening but was able to take a few images up high above it.
11/10/2021
Yong Sheng
Dieter Hartmann
CT griz
9
0
Equipment Failure
There is a problem with the limit switch on the dropout. Brian said that he will send someone to check it out.
11/03/2021
Tomomi Otani
Johnathan Hodge
Andor CCD
10
0
No problems to report
Hot Subdwarf photometry observation for time series. Everything was working well tonight.
11/02/2021
William Keel
2 cats
Andor CCD #2
3
5
Clouds
NGC 1531, dodging clouds. Focus 128300, guider focus 8400. Very poor seeing early in night.
11/01/2021
Tomomi Otani
none
Andor CCD
7
0
Clouds
Hot subdwarf photometry time series. When I slew the telescope, the RA did not move. If I click the "slew" button again, it started moving with no problem.
10/30/2021
Tomomi Otani
Johnathan Hodge
Andor CCD
8
2
Equipment Failure
When we logged into the telescope control screen at the beginning of the night, we saw DEC limit (we saw all "N", "S", "W", "E") and could not slew the telescope. We also could not slew telescope using the pad. However, the airmass was 1.4. Peter helped me to fixed it. Peter said that it happened due to Loose wire on the telescope, or faulty tilt sensor, or something like that.
10/29/2021
Paulina Karczmarek
None
Andor CCD
7
1
Clouds
Thin clouds for majority of the night, so backup targets observed. Thick clouds in the last 2h of the night hindered observations of even backup targets.
10/28/2021
Tomomi Otani
Johnathan Hodge
Andor CCD
10
0
No problems to report
Hot subdwarf photometry observation. No problems to report.
10/27/2021
Tom Rutherford
None
CCD
8
1
Software/Control PC Issues
Standard stars, variable star. No issues other than one I caused by closing the wrong window-- a full restart fixed it.
10/26/2021
XIANMING HAN
none
CCD
7.5
0
Other
eclipsing binary. The telescope was parked beyond the west limit from previous day, and I could not move it back. Peter from ACE helped moving the telescope back to operation range. The CCD camera was also way out of focus such that no stars were visible initially and got me confused for awhile, but I was able to realize the problem and fixed the focus.
10/26/2021
Tomomi Otani
None
Andor CCD
10
0
No problems to report
Hot Subdwarf photometry time series. No problem to report. Focus 129000 (no filter). CCD temperature: -55 C
10/24/2021
Tom Rutherford
None
CCD
6
2
Software/Control PC Issues
Exoplanet. Image downloads were VERY slow (several minutes for each bias frame). Rebooted the CCD computer and it worked after that. The CCD sequence control on the telescope computer failed to work so rebooted the telescope computer as well-- took several tries to get it back up and running.
10/23/2021
Todd Hillwig
None
Andor CCD
8
0
No problems to report
Testing new Andor CCD. PN central stars.
10/22/2021
Keefe Kamp
Gokuldass, Norton
CCD
9
0
No problems to report
TESS Follow-up Observations. Was able to connect this time with Ubuntu. No new problems to report
10/21/2021
Tom Rutherford
None
CCD
8
0
No problems to report
Exoplanet Transit, variable star
10/20/2021
Keefe Kamp
Gokuldass, Norton
CCD
9
1
Software/Control PC Issues
TESS follow-up observations. Could not connect to VPN from Ubuntu but from Mac worked. Ace froze when abort was pressed. Simple closing and reopening ACE fixed it. Weird banding artifact present in observations strongest in B caused low SNR.
10/19/2021
Tom Rutherford
None
CCD
8
0
No problems to report
No problems-- all went well.
10/18/2021
Gergely Hajdu
Paulina Karczmarek
Andor camera
8
0
No problems to report
Observing program: absolute photometry in Sloan bands for local comparison stars of bright variables (Cepheids, RR Lyrae). Problems: sometimes offsets as large as RA -60, DEC -150 were necessary to move the fields to the desired position.
10/17/2021
Keefe Kamp
Gokuldass
CCD
9
0
No problems to report
TESS Follow up Observations of B stars and High Proper motion stars
10/15/2021
William Keel
AY203 class, 2 cats
Andor CCD #2
10
0
No problems to report
Class nebulae, cool galaxies. Seeing very good 2nd half of night. Focus 126300 (1000 higher for [O III\ 5007 filter). Guider focus 9900. Telescope did bob N-S around declinations -10 to -20.
10/14/2021
Michael Fauerbach
NA
CCD
8
2
Clouds
Clouds started rolling in towards moring.
10/13/2021
William Keel
2 cats, Astro 203 lab class
New Andor CCD
10
0
No problems to report
Class demo nebulae, galaxies in BVR and H-alpha. Focus 125700, 5007 filter focus 1000 smaller. Guide focus 8400. Stray light is still there, with guider in extreme positions as well as center. Some targets took two presses of the GO button to complete a slew.
10/13/2021
Carlos Zuluaga
S. Levine
Andor Camera
3
0
No problems to report
Outer Solar System astrometry.
10/12/2021
Ted von Hippel
Giovanni Bacon, Jack Wood
imaging
4
0
No problems to report
Student observing for a class. They needed only the first 1/2 of the night and when I put out a call for users for the 2nd 1/2 of the night (admittedly with very little notice), there were no takers.
10/09/2021
Martha Leake
Logan Spooner
ARC camera
10.25
0
Other
Student projects, asteroid lightcurves and studies in parallax. Weather generally good. Images occasionally oscillated up and down close to transit. Confused by start time. Some oscillation of tracking errors during multiple exposures.
10/08/2021
Ken Rumstay
None
CCD camera
1
9
Humidity
AGN monitoring program. Everything worked fine for about an hour, long enough to get bias, flat, and dark images. When slewing to my first target, the telescope stopped, and the TOS indicated that it had reached the S and W limits. It was actually pointed northeast of the zenith. Peter was able to get the telescope moving again, but by then clouds had rolled in and the humidity climbed to over 80%. I gave up after midnight EDT. Previous observers reported similar limit issues in August. Observers should note that the problem may recur.
10/06/2021
Michael Fauerbach
NA
CCD
10
0
No problems to report
Asteroid phptometry
10/05/2021
Ken Watanabe
Nicolas RImbert
CCD
5
0
No problems to report
Stars tend to have triangle shape. Blazars.
10/03/2021
Stephen Levine
C. Zuluaga, A. Bosh
Andor Camera
4
0
No problems to report
Outer solar system astrometry. No issues. The Andor camera worked well.
10/02/2021
XIANMING HAN
Sera Shaughnessy
CCD
7
3
Equipment Failure
eclipsing binary and asteroid Camera computer was not reachable initially. ACE (and possibly others) helped recover the computer.
10/01/2021
XIANMING HAN
none
CCD
4
0
No problems to report
eclipsing binary, sharing the night
10/01/2021
XIANMING HAN
none
CCD
4.5
0
No problems to report
eclipsing binary, sharing the night
10/01/2021
XIANMING HAN
none
CCD
5
5
Clouds
eclipsing binary
09/30/2021
Walter Van Hamme
Samec
24" scope, CCD
0
11
Clouds
Observations and analysis of eclipsing binaries.
09/30/2021
Walter Van Hamme
samec
24", CCD
10
0
Clouds
Light curve observations and analysis of GSC75170134. I complete curve single coverage.
09/28/2021
Todd Hillwig
None
Andor CCD
6
0
No problems to report
PN central stars.
09/27/2021
Keefe Kamp
Gokuldass
CCD
6
4
Software/Control PC Issues
B star observations with CCD. Data stopped being saved properly because the disk became full.
09/26/2021
Paulina Karczmarek
none
Andor CCD
10
0.5
Software/Control PC Issues
"Classical Cepheids in the SDSS photometric system". The directory with data is virtually full, and that means that new exposures are not being saved, instead a 1KB file is created with no content. Deletion of a directory with our previously collected data solved the problem for now.
09/23/2021
Judi Provencal
none
Andor
5
0
No problems to report
Imaging for observational astronomy class
09/21/2021
Tom Rutherford
None
CCD
6
0
No problems to report
Asteroid. No issues at all.
09/19/2021
Paulina Karczmarek
Gergely Hajdu
Andor CCD
10
0
Clouds
"Classical Cepheids in the SDSS photometric system", thin clouds the whole night.
09/19/2021
Paulina Karczmarek
none
Andor CCD
10.5
0
No problems to report
"Classical Cepheids in the SDSS photometric system"
09/19/2021
Paulina Karczmarek
none
Andor CCD
5.5
4.5
Clouds
"Classical Cepheids in the SDSS photometric system", clouds starting the middle of the night prevented further observations.
09/15/2021
Todd Hillwig
None
Andor CCD
6
0
No problems to report
PN central stars
09/15/2021
Todd Hillwig
None
Andor CCD
6
0
No problems to report
central stars of planetary nebulae.
09/13/2021
Michael Fauerbach
NA
CCD
10.5
0
No problems to report
Asteroid photometry. Some thin clouds, else no issues.
09/12/2021
Martha Leake
Logan Spooner and Cera Purdy
ARC camera
1
8
Humidity
Asteroid photometry and other objects for Freshman Research Experience. Never opened. Clouds and high humidity issues early on; humidity predicted to persist until morning twilight. CTIO weather page and Clear Outside data mostly agreed: fog, possible snow and frost, humidity over 75% to 90%. Took just biases and darks.
09/11/2021
C. Zuluaga
S. Levine
Andor Camera
0
11
Clouds
Outer solar system astrometry.
09/03/2021
Michael Fauerbach
NA
CCD
8.5
2
Equipment Failure
Asteroid photometry. The night was clear and the telescope was working great for most of the time. Howver, there were a few of my targets were the telescope would not track. Switched tracking off and on helped a bit for one frame, and then it would streak again. All targets were well above 40 degrees elevation. Moved to a different target and all worked perfectly again.
08/31/2021
William Keel
3 cats
Andor CCD
10
0
No problems to report
Galaxies. Telescope oscillated rapidly N-S when north of zenith, even in 1-second exposures. Following an object for long times improved but did not eliminate this. Focus 126800, guider focus 500.
08/30/2021
Saida Caballero
Kamp, Malagon
CCD
8
0
No problems to report
B star monitoring
08/29/2021
Tanner Norton
Caballero, Van Putten, Gokuldass, Malagon
CCD
6
0
No problems to report
Be Binary Survey
08/28/2021
Todd Hillwig
None
Andor CCD
5
5
Clouds
PN central stars. Thin clouds throughout the first half of the night, leading to overcast halfway through.
08/21/2021
Núria Torres-Albà
Dieter
SARA CT
10.5
0
No problems to report
Clemson University Observations of blazars. No problems to report.
08/20/2021
Robert Berrington
none
ARC Camera
0
10
Clouds
short period variable stars
08/19/2021
XIANMING HAN
none
CCD
0
10
Clouds
none
08/18/2021
Yong Sheng
Dieter Hartmann
griz
7.5
2.5
Software/Control PC Issues
The telescope was stuck at the RA and Dec limits
08/16/2021
XIANMING HAN
none
CCD
0
10
Clouds
The telescope was stuck at the east limit in addition to clouds. Emailed ACE to report this problem.
08/16/2021
XIANMING HAN
none
CCD
10
0
No problems to report
eclipsing binaries, jupiter
08/11/2021
Michael Fauerbach
NA
CCD
10
0
No problems to report
Asteroid photometry
08/10/2021
Steve Gillam
Kiara Richardson & Chloe Nelson
Andor CCD
13
0
No problems to report
ROAp stars.
08/08/2021
Stephen Gillam
None
Andor CCD
3
7
Humidity
Clouds, humidity, and low temperatures forced closedown.
08/05/2021
Stephen Levine
C. Zuluaga, A.Bosh
Andor Camera
4
0
No problems to report
Preparatory observations for occultation tomorrow.
08/04/2021
Ken Rumstay
Max tha cat
CCD camera
9
0
No problems to report
Observed active galaxies. The telescope systematically pointed about three to four arcminutes north of target. About 30-60 minutes east of the meridian (depending upon the declination) the telescope appeared to oscillate in declination, resulting in images severely elongated in the north-south direction. The connection to the CTIO computers was terminated on three occasions during the night as noted by earlier observers), but reconnecting posed no problem.
08/02/2021
Martha Leake
none
Andor
12
0.5
Connection Issues
Asteroid photometry. Some connection issues getting started; used the full reconnection routine (and ace-gui message appeared). Don't forget Autosave. Can't see sky at CTIO because they are possibly disconnected? Two slews needed another "Go" or time to finish. One radmin drop of Sara Telescope. Last observations at altitude 53 to east resulted in trails first tilted right then left, then straighter (N was down, EW horizontal) Pointing good. Maxim timed out and was not used. Seeing about 1"and weather great.
07/31/2021
Keefe Kamp
1 Chinchilla
CCD
11
0
Connection Issues
Be star observations. Had a problem where VPN kicked me out and wouldn't let me back on. A restart of my machine fixed that. Camera computer also alerted that C: Drive was low on memory.
07/30/2021
Alex McDaniel
Dieter
Andor CCD
12
0
No problems to report
Blazar Photo z measurements. No issues tonight
07/28/2021
Alex McDaniel
Dieter Hartmann
Andor CCD
12
0
No problems to report
Blazar Photo z measurements. No issues tonight
07/19/2021
Van Hamme
Samec
24"
12
0
No problems to report
Observations of interacting binaries, V2716 Oph and GSC 7517 0234. Continued Observations. A few clouds at the start only.
07/19/2021
Van Hamme
Samec
24"
null
null
No problems to report
Observations of interacting binaries, V2716 Oph and GSC 7517 0234. Began Observations.
07/17/2021
Michael Fauerbach
NA
CCD
11
0
No problems to report
Telescope computer was unresponsive a couple of times -looked like a lost connection. Only the telescope computer had these issues. Blanco and SOAR webcams stopped updating images around 11:30pm local time.
07/15/2021
Harum Ahmed
Kent Montgomery
Andor
8.5
0
No problems to report
Asteroid Photometry Ended observation due to reaching 30 degrees altitude. Weather: Clear skies.
07/14/2021
Harum Ahmed
Kent Montgomery
Andor
9
0
No problems to report
Asteroid Photometry Weather: Clear skies.
07/13/2021
James Webb
None
CCD
0
10
Connection Issues
My internet flaked out early and never had a chance to observe.
07/13/2021
Harum Ahmed
Kent Montgomery
Andor
9
0
Other
Asteroid Photometry Problems: The telescope moved North for 3 min at Alt. of 32 degrees, H.A. was -4:46:41 and Azimuth was 126. Had streaks in the image. Weather: Clear skies.
07/12/2021
Harum Ahmed
Kent Montgomery
Andor
7
1
Other
Asteroid Photometry Problems: Closed observation due to reaching the limit of 30 degrees altitude. Weather: Clear with thin clouds.
07/10/2021
Stephen Gillam
None
Andor camera
9
1
Other
rOap stars. Stormy weather in Daytona Fl. caused internet outage which prevented connection to SARA-CT at the start of the night.
07/08/2021
William Keel
3 cats
Andor CCD
11
0
No problems to report
Dead quasar survey, H II regions. Focus 128200 start, 126700 end, guider focus 0. Some equatorial targets took multiple slew commands to get to the right place.
07/07/2021
XIANMING HAN
none
CCD
10
0
No problems to report
binaries, Jupiter satellites
07/07/2021
Ross Silver
Dieter, Yong
CCD
0
10
Equipment Failure
Photometric redshift of blazars. The DEC tracking would not work. We reset the whole system and the tracking but could not get it to work.
07/03/2021
Mabel Clausen
Kurtis Williams
Andor
0
9
Clouds
Thick clouds at the start of the night unlikely to clear.
07/03/2021
Mabel Clausen
Harum Ahmed
Andor
5
4
Clouds
White Dwarf Time Series Photometry Problem: Telescope moved South for Image 98. For more details see last nights report was the same. Weather: Thickening clouds had to close early
07/02/2021
Mabel Clausen
Harum Ahmed & Kurtis Williams
Andor
9
0.5
Equipment Failure
White Dwarf Time Series Photometry Problem: Near start of night and at end of target run telescope moved north/south over several exposures HA: -2:51, Az: 84, ALT: 50 Telescope moved south 87 arc sec over 9 min HA: +3:39, Az: 269, ALT: 39 Telescope moved north 79 arc sec over 15 min We were not using auto guider Telescope pointing was steady otherwise Note: ACE disabled telescope limits: at their request we kept telescope at altitude greater than 30 degrees all night Weather: High thin Cirrus all night Seen: 2-3 arc sec Focus: 12770
07/01/2021
Gergely Hajdu
Paulina Karczmarek
Andor camera
10
2
Equipment Failure
Observing program: backup program (griz time series of V445 Oph and V Ind), due to thin clouds making the original program unfeasible (absolute photometry of Cepheid comparison stars in the Sloan bands). There were technical problems on two occasions, first when observing V445 Oph around azimuth ~ 0, the camera FoV would produce very strange vertical (DEC direction) PSFs while exposing. Sometimes the length of these reached over 300 pixels! This lasted for about an hour, ending around UT03. Secondly, while observing V Ind, around UT06:50 the telescope got completely stuck. I couldn't move it with giving it movement orders, offsets, or the paddle. The only thing I could do is to close the mirror cover and the dome, to leave the telescope in a safe state. As this happened near the end of the night, we didn't lose too much time, but I couldn't attempt doing sky flats.
06/27/2021
Stephen Gillam
None
Andor Camera
13
0
No problems to report
R0Ap stars.
06/25/2021
Michael Fauerbach
NA
CCD
12
0
No problems to report
no issues, other than the Full Moon
06/23/2021
Gary Henson
Rutherford
Andor CCD
7
2.5
Clouds
Asteroid 18488 Light Curve Dec Oscillation at around -1 HA while tracking object. Problem appeared as reported by other observers, went to Zenith park and then back on target with no further occurrences. Band of cirrus clouds after midnight. Seeing and tracking with autoguider excellent most of night.
06/20/2021
Tomomi Otani
Andy Baran
Andor CCD
12
0.3
Software/Control PC Issues
Photometry time series of a hot subdwarf star. Thank you very much, Brian, for adding the decimal points to the observation time of the image header. Everything was good but the telescope started moving to the DEC direction at the Hour Angle ~ +3 hours. This was the same issue that we experienced the last night, so it may repeat again in future.
06/19/2021
Otani
Andy Baran
Andor CCD
12
0.5
Software/Control PC Issues
Photometry time series to check pulsations. The telescope suddenly started moving to the DEC direction (the target's y position moved 2 pixels/image). So I moved the telescope to the zenith and moved it back to the target, then the issue was solved. The camera connection suddenly disappeared while taking images, so I used the regular procedure to reconnect the camera. The header's observation time did not have decimal digit (so we could not obtain less than 1 second accuracy).
06/18/2021
Michael Fauerbach
NA
CCD
9
3
Clouds
Clouds off and on throughout the night. Heavy cloud cover in the morning.
06/16/2021
Kurtis Williams
Mabel Clausen & Harum Ahmed (REU students)
N/A
0
12
Clouds
A text note was open on the desktop saying that due to testing the telescope limits had been disabled and the telescope should not be used. Was that a one-day affair, or is this ongoing for multiple nights? If the latter, perhaps a note should go out to sara-tech... (This didn't affect us since clouds were thick and the weather forecast poor)
06/14/2021
Stephen Levine
C. Zuluaga, A. Bosh
Andor Camera
1
9
Equipment Failure
Outer solar system astrometry and Pluto occultation. When trying to pre-set to the night's occultation field, which was going to be very low and rising, the telescope became stuck in the limits and would not back out. No motions responded from the hand paddle. In the end, we closed things up as it was.
06/13/2021
Yong Sheng
None
CT griz
13
0
No problems to report
Clemson Blazar study.
06/11/2021
Ken Rumstay
None
CCD camera
12.5
0
No problems to report
AGN monitoring program. An excellent night, with a dark sky and sub-arcsecond seeing.
06/10/2021
William Keel
3 cats
Andor CCD
7.5
0
No problems to report
Nebulae, galaxies behind dust. Focus=126500, guider focus 9000.
06/09/2021
Stephen Levine
C. Zuluaga, A. Bosh
Andor Camera
2
0
No problems to report
Outer solar system astrometry.
06/07/2021
Steve Gillam
None
Andor Camera
12
0
No problems to report
ROAp stars
06/02/2021
Michael Fauerbach
NA
CCD
7
4
Equipment Failure
As mantioned in my email to sara-tech, the whole telescope computerACE crashed after a CCD abort. After recovering, the Default DEC Encoder fault kept flashing. Tried to reboot computer etc but nothing got rid of it. I was able to park, close the covers and close the dome shutters -at least it said it did. I have not seen this error before, and I was not able to recover from it.
06/01/2021
XIANMING HAN
none
CCD
10
0
No problems to report
eclipsing binaries
06/01/2021
XIANMING HAN
none
CCD
0
10
Clouds
none
06/01/2021
XIANMING HAN
none
CCD
0.25
10
Clouds
eclipsing binary
05/31/2021
Tom Rutherford
None
CCD
7
0
No problems to report
Landolt Fields. Open Clusters. No problems at all.
05/28/2021
Tom Rutherford
None
CCD
0
9
Clouds
Clouds all night-- unable to open.
05/27/2021
Tom Rutherford
None
CCD
7
0
No problems to report
Landolt fields. No problems-- everything worked perfectly.
05/16/2021
Fauerbach
NA
CCD
8
3
Clouds
asteroid photometry. Clouds the first three hours, else fine
05/15/2021
Michael Fauerbach
NA
CCD
8
2
Clouds
Clouds early in the morning Filterwheel 2 is not responding to commands
05/14/2021
Stephen Levine
C. Zuluaga, A. Bosh
Andor Camera
11
0
No problems to report
Outer solar system astrometry. One instance where the tracking stopped. Not clear why. Able to re-enable and keep going.
05/13/2021
Leila Alamos
Stephen Gillam
Andor Camera
12
8
Clouds
Observing Program: ROAp stars Problems: It was very cloudy during the entire observation Suggestion: Filter wheel 2 might indicate a problem due to its constant blinking. *relates to the May 11th and May 12th night of observation
05/13/2021
Leila Alamos
Stephen Gillam
Andor Camera
12
6
Humidity
Observing Program: ROAp stars Problems: Humidity was very high towards the end of the night and so we had to pack up early. Suggestion: Filter wheel 2 might indicate a problem due to its constant blinking.
05/10/2021
Stephen Levine
C. Zuluaga, A. Bosh
Andor Camera
0
11
Clouds
Outer solar system astrometry.
05/09/2021
Steve Gillam
None
None
0
13
Connection Issues
ROAp stars A recent update to windows 10 killed the Cisco VPN client on my computer. I was unable to reinstall it. This is the 7th time in the last 3 years that I have had to reinstall the Global VPN client, the Cisco VPN client, the DNE program, and run Winfix.exe after a Windows update. Running winfix.exe corrupted the Windows Installer neccessitating a complete reinstallation of my Operating system. I strongly recommend replacing the Cisco VPN service, which is no longer supported by Cisco, with one that is supported and compatible with Windows 10.
05/07/2021
Michael Fauerbach
NA
CCD
0
10
Clouds
Never opened cloudy all night
05/07/2021
Todd Christopher Hillwig
None
Andor CCD
10
0
No problems to report
None.
05/01/2021
Todd Christopher Hillwig
None
Andor CCD
10
0
No problems to report
PN central star monitoring
04/29/2021
Tom Rutherford
Henson
CCD
5
0
No problems to report
Southern T Tauri stars. No problems-- everything worked well.
04/27/2021
Gary Henson
Rutherford
Andor CCD
4
1
Equipment Failure
Lost first hour when Altitude Limit tripped. Telescope not near horizon, but Slew not recovered until ACE turned off switch to move telescope. Pointing, tracking, filter wheel sequence all worked fine so observations excellent once started.
04/26/2021
Walter Van Hamme
samec
camera
null
0
No problems to report
Observations of active solar type binaries.
04/24/2021
William Keel
3 cats
Andor CCD
11
0
No problems to report
H II regions near and far. Explored stray light mitigation by doing images with shutter open and telescope covers closed. Early on, a span of panic when it appeared that stars left charge trails along -y but this seems to have been asymmetric telescope bounce. Focus 126800, guide focus 2000. Pleased to see that the Andor CCD can recover control after an abort just by reopening the "observations" panel in ACE software. Guider was not able to correct for NS telescope motions near -65, 1hr west.
04/22/2021
William Keel
3 cats
Andor CCD
5
5
Clouds
Galaxies in H-alpha, checking out stray light. Focus=127800, guider focus 2000. Guider parameters had been corrupted; reset to X speed = Y speed=1, angle=0. The stray light between filters is puzzling - it looks brighter (in ADU/minute) in H-alpha than in R, and only slightly dimmer in B than in R. Its contrast across the field does seem to change with filter.
04/21/2021
Alex McDaniel
Yong Sheng, Dieter Hartmann
CT
6
5
Humidity
Blazar Photo z measurements
04/18/2021
Saida Caballero
Kamp, Le, Fowler, Richards, Miorelli
CCD
12
0
No problems to report
Photometric survey Be stars.
04/17/2021
Todd Hillwig
None.
Andor CCD
9
1
Clouds
PN central star photometry
04/16/2021
William Keel
3 cats
Andor CCD
11
0
No problems to report
Dead quasar survey, H II regions, stray-light checkout. Focus=128300, seeming finicky tonight.
04/15/2021
Stephen Levine
C. Zuluaga
Andor Camera
11
0
No problems to report
Outer solar system astrometry. No issues to report.
04/14/2021
Stephen Levine
C. Zuluaga
Andor Camera
12
0
No problems to report
Outer solar system astrometry, and basic testing of new Andor camera to understand the performance in the various modes. Looks good.
04/12/2021
Ken Rumstay
None
CCD camera
9
0
No problems to report
AGN monitoring program No problems to report Optimum focus = 127800 at T = 16 Celsius Telescope pointing after slew not perfect, but target was always within CCD filed.
04/11/2021
Ross Silver
Dieter, Yong
CCD
11
0
No problems to report
Photometric redshifts of BL Lacs.
04/06/2021
XIANMING HAN
Andrew Gebken
CCD
8.5
0
No problems to report
eclipsing binary
04/06/2021
XIANMING HAN
Andrew Gebken
CCD
8
0.5
Clouds
eclipsing binary
04/06/2021
XIANMING HAN
Andrew Gebken
CCD
8.5
0
No problems to report
eclipsing binary
04/06/2021
XIANMING HAN
Andrew Gebken
CCD
8.5
0
No problems to report
Eclipsing binary
04/06/2021
William Keel
3 cats
Andor CCD
11
0
No problems to report
Dead quasar survey, Andor CCD checkout. Focus 128000, guider focus 2000. Found potential gotcha when recovering from sequence abort - CCD came back in 3 MHz readout instead of previously set 1 MHz; bought myself some noise for the next hour or so. There is a diffuse pattern in darks at about 1 ADU/minute which looks like a light leak but persists at night so it seems to be an actual property of the chip.
04/01/2021
Todd Hillwig
None
Andor CCD
10
0.5
Equipment Failure
Telescope routinely pointed about 10 arcminutes to the north. This was corrected every time by simply hitting "Send" a second time, at which point the telescope pointed correctly with the target in the inner third of the chip.
03/29/2021
Gary Henson
Rutherford
Andor CCD
0
6
Other
NOTE: this happened night of 4/28 so 4/28 report is really for 4/26, don't see way to edit old reports so submitting with wrong date for some completeness of logs. Power Failure at Dome just after flats, Power out 4+ hours but all systems came back up once power restored. Uncertain of failure cause and night half gone so closed up and quit early.
03/28/2021
Tom Rutherford
Henson
CCD
6
0
Other
Asteroid light curve. Due to the filter wheel issue, I set wheel #2 to clear (where it was supposed to have been already, but wasn't) and moved wheel #1 to "R." Both wheels seemed to behave normally. Did no filter changes from this point onward-- everything else behaved properly and I had no problems with the scope except for a sudden jump toward the south by an arc minute or so (something that has happened to me before with CT) halfway through the run.
03/24/2021
James Webb
keslo Estil
CCD
4
0
Other
Imaging nebula and galaxies Tracking was horrible at 2 hours before transit. Not even the auto guider could correct guiding errors. Every where else, tracking was great.
03/22/2021
Yong Sheng
Alex McDaniel, Dieter Hartmann
g, r, i, z filters of CT
1
10
Equipment Failure
We can't observe because of the filter issue on the wheel 2. It won't change to the correct filter, and it got stuck.
03/18/2021
Stephen Levine
Carlos Zuluaga
Imaging camera
11
0
No problems to report
Outer solar system astrometry. At sunset, clear, T=12.6C, RH=56%, Windspeed=4.4mph A smooth night.
03/17/2021
Keefe Kamp
Caballero, Zamora
CCD
10
0
No problems to report
Observed Comets for ObsAstro Class and Standard Feilds with help from guide camera for ETC. Everything behaved normally
03/13/2021
Todd Christopher Hillwig
None
FLI CCD
10
0
No problems to report
PN central star monitoring. Only used Filter Wheel #1 and had no issues with it.
03/12/2021
William Keel
3 cats
FLI CCD
10
0
No problems to report
Dead quasar survey, galaxy dust lanes. Focus 132200. guider focus 2000.
03/11/2021
Martha Leake
students
FLI Cam
2
7
Clouds
Added comment: I attempted to change filters from SDSS_R to empty, in order to use the Bessel filters in Wheel 1. The filter change didn't progress, and looked like a blockage of the frame. Ran the initialization of Wheel 2 with unknown success. Filter remains sdss_r.
03/11/2021
Martha Leake
Students in Tools class
FLI Cam
2
7
Clouds
80 percent high clouds, star fields visible through it, however. Predicted high humidity reached near morning twilight. Student participation delayed because of local connection issues.
03/08/2021
Ross Silver
Yong, Alex, Dieter
CCD
0
10
Equipment Failure
Photometric redshift of blazars. Filter wheel 2 is still not reliable and working consistently, therefore we could not observe.
03/06/2021
Amy Lovell
CCD
10
0
No problems to report
Objects for lab - images and lightcurves
03/05/2021
Saida Caballero
Malagon
CCD
10
0
No problems to report
Standard star fields. No issues to report.
03/04/2021
Amy Lovell
AST200L: Kimbarl King Krahm Nguyen
CCD
8
0
No problems to report
Imaging a variety of objects with AST200 lab students
03/03/2021
Keefe Kamp
Caballero
CCD
4
0
No problems to report
Observed Various Targets with Observational Astronomy Class. Telescope tracking seemed slightly off as stars were elongating even with short (~10 sec) exposures.
02/26/2021
Saida Caballero
N/A
CCD
0
10
Clouds
Too cloudy to open. From previous night (24 Feb) forgot to mention that the guide camera was not working. Did not test it tonight.
02/25/2021
Saida Caballero
Kamp, Klare, Le, Zamora
CCD
10
0
No problems to report
Class project for creating an Exposure time calculator. No problems to report. CCD was back online after having stopped when aborted the night before.
02/24/2021
Saida Caballero
Kamp
CCD
10
0
Software/Control PC Issues
Starting taking darks before the end of the night and had to abort the observations for morning flats. Followed the restart instructions for the camera (even restarted both computers) and was not able to get the camera back online. Message is "Exposure start failed."
02/21/2021
Amy J Lovell
CCD
6
5.5
Equipment Failure
Started okay, but then camera disconnected, required ACE intervention. Once restored, no problems to report.
02/20/2021
Núria Torres-Albà
Yong, Alex, Dieter
SARA CT
7
1
Software/Control PC Issues
Clemson University blazar observations Problems: Sequence tool got stuck and we had to restart ace control software
02/18/2021
Amy Lovell
Kimbarl, King, Krahm, Nguyen
CCD
9.5
0
No problems to report
A few of the images had some streaks of interrupted tracking, otherwise all was well.
02/16/2021
Martha Leake
Students in two classes observed
FLI Cam
10.5
0.3
Equipment Failure
Demonstration objects and asteroids (light curves). Most everything worked well, but a few glitches. On last asteroid in the 12:54 RA, -9 17 Dec (meridian area), a 240 second exposure had short streaks--a slippage? This short streak was repeated on each subsequent frame--not as if tracking stopped--it was engaged (and not red), but perhaps a slip. Two slews were so long I used the Stop button, then went to the object without problem. One of these was adjusting HA to 00:00 for the final parking at +15 deg Dec. After the streaks I moved to zenith park, tested a no-tracking streak of different lengths, and then tracked (aimed near zenith) and tracked the area. No problems appeared. I expected high humidity tonight but it did not exceed 72% (at start). Early on, noted ds9 wasn't displaying data (just image); Tried the XPA. Closed ds9, started Ace connector on camera computer, followed by reopening ds9 and using the telescope computer to take pictures. This worked (and is in the manual). Focusing poor, seeing not the 0.6 " advertised on Tololo conditions. Remembered to disconnect vpn link when done.
02/14/2021
Ross Silver
Dieter Hartmann
CCD
0
9
Equipment Failure
Photometric redshifts of BL Lacs.
02/13/2021
Martha Leake
Logan Spooner
FLI Cam10.7
10.7
0
No problems to report
Student targets (M92, LMC) and asteroids 445 Edna and 13093 Wolfgangpauli. Parked telescope at HA = 00:00 and Dec = +15. NIce weather. No problems other than busy day just prior to run.
02/11/2021
William Keel
3 cats
FLI CCD
8.5
0
No problems to report
Merging and backlit galaxies. Focus 132500, guider focus 0. Used only filter wheel 1 with no problems. When moving to park at dec +30, ACE control gave error message about faulty dec encoder. Compared to KP, CT computer gave UT 22 seconds slow (and KP agreed with cellphone time to a second or so).
02/10/2021
James Raymond Webb
Viviana Arroyave, Keslo Estil
CCD
6
0
No problems to report
Blazar monitoring and imaging galaxies
02/08/2021
Yong Sheng
Meenakshi Rajagopal
SARA CT
0
10
Software/Control PC Issues
Study BL Lacs Problem: Can't use the filters in wheel 2, the filter options are in grey
02/07/2021
Ken Rumstay
None
CCD camera
0
10
Equipment Failure
The #1 filter wheel is not operating, as noted by Amy Lovell last night.
02/06/2021
Amy Lovell
none
CCD
2.5
8
Equipment Failure
Filter Wheel 1 is unresponsive: INIT Wheel 1 just continually displays "Finding HOME for Wheel 1" and never returns.  All filters in Wheel 1 are grey.  The filter wheel window also persists after closing the Science Camera.  Any attempt to fool the wheel by selecting a filter in the sequence tool is similarly met with no response. The camera reset procedure (from the text file on the desktop) does not fix this, whether with the Observations window or the whole ACE software. Reboot of either or both computers in any order doesn't fix. After the resets, etc, the camera will expose and write an image, with "---" in the FILTER field, but the filter may have been stuck on B or between B and V.
02/05/2021
William Keel
3 cats
FLI CCD
8.5
0
No problems to report
Focus 133000; Guider focus 0. Near the meridian at dec -69, there were frequent instances of the telescope bobbing N-S by 5” or so that the guider took 5 minutes to correct. After recent nights, I kept closer track of the CD temperature and took multiple darks during the night to see whether hot pixels are really changing that fast. Two darks during twilight were suspiciously bright - light may be leaking around the shutter.
02/04/2021
Walter Van Hamme
Samec
FLI
8
0
No problems to report
All went well. Observations on EI Oct and WISEJ051352.5-170113. The network did kick me out 20 or so times but I was able to easily connect again.
02/03/2021
Saida Caballero
Kamp, Class
CCD
10
0
No problems to report
Class observing and standard star field. X-stage guider did not seem to like moving to the W fields. Had to reinitialize. Guiding worked very well. We did reach the horizon limit when trying to slew away from it, but were able to move the telescope with the paddle and slew normally.
02/02/2021
Keefe Kamo
Saida Caballero, Natalia Malagon
CCD
11
0
No problems to report
Observed Standard Fields for CCD calibration utilizing Guider
02/01/2021
William Keel
3 cats
FLI CCD
8
0
No problems to report
Galaxies, supernovae, Surveyor 2 Centaur stage. Focus 131500, guider focus 0. Guider stage will only reach positions 0 and E1. There was an error message FAULTY ABSOLUTE DEC ENCODER on ACE screen; unsure when it appeared.
01/31/2021
Keefe Kamp
Caballero, Le, Klare
CCD
11
0
No problems to report
Standard Star Feilds taken for calibration. The only problem that occurred was accidentally pressing abort twice and needing to follow the notepad instructions to reconnect the camera.
01/31/2021
Amy Lovell
King
CCD
3
3
Connection Issues
student training, images for lab analysis, connection problems at this end, tried to find the right pointing offset a few times
01/31/2021
Amy Lovell
Krahm
CCD
2
7
Clouds
Student training, thin clouds rolled over after not too long
01/27/2021
Van Hamme
Samec
FLI camera
9
3
Equipment Failure
Observations of Eclipsing Variables, EI Oct and WISEJ051352.5-170113. Network problems. connection very weak. Closed down multiple times. Very slow reacting.
01/27/2021
Walter Van Hamme
Samec
FLI camera
12
0
No problems to report
Observations of Eclipsing Variables, EI Oct and WISEJ051352.5-170113.
01/27/2021
Van Hmme
Samec
FLI Camera
12
0
No problems to report
Observations of Eclipsing Variables, EI Oct and WISEJ051352.5-170113.
01/25/2021
Tomomi Otani
Andy Baran
CCD
9
0
No problems to report
Hot Subdwarf stars photometry time series. Focus 134700 at the beginning of the night, and I needed to increase it later. The weather was good but some clouds came during the second half of the night. Telescope was working well.
01/23/2021
Tom Rutherford
Gary Henson
CCD
5
3
Clouds
Cloudy first part of the night, but cleared around local midnight. The telescope and CCD worked perfectly.
01/22/2021
Walter Van Hamme
Samec
FLI Camera
8
0
No problems to report
Observations of Eclipsing variables WISE051352,5-170113 and EI Oct.
01/22/2021
Walter Van Hamme
Samec
FLI Camera
8
0
No problems to report
Observations of Eclipsing variables WISE051352,5-170113 and EI Oct.
01/20/2021
Tomomi Otani
No one
CCD
7
0
No problems to report
Subdwarf B star photometry time series for searching for companions. Everything was working well and the weather was good too.
01/19/2021
James Webb
Keslo Estil
CCD
6
0
No problems to report
Galaxy imaging
01/18/2021
Núria Torres-Albà
Dieter Hartmann
-
8
0
No problems to report
Clemson University Blazar Observations. No problems to report.
01/15/2021
Ken Watanabe
Nick Rimbert
None
0
9
Connection Issues
Not able to keep connection using cisco vpn on win 10. Not other way to connect?.
01/09/2021
Ken Rumstay
None
CCD camera
9
0
No problems to report
AGN monitoring program No problems to report. Telescope tends to point north of target, as noted by Bill Keel. Optimum focus 131000 at 12 Celsius. CCD temperature was -28 Celsius at start; it cooled to -30 within an hour and remained there.
01/08/2021
William Keel
3 cats
FLI CCD
8.5
0
No problems to report
Test drive, galaxies in H-alpha. Focus=133000. Even after initialization, the guider x-stage has limited movement; to the west it can’t quite reach W1. The NSEW directions are 180 degrees rotated from their previous directions; +x is to the east and +y to the north. The guider settings for x and y speed and angle are 1,1,0 (sign changes on the speeds from previous values). The guider focus is close to 0. Guiding works well; in some parts of the sky there were oscillations of 1-2” lasting several minutes, as if the slow-motion controls didn’t have enough authority in those locations. There was one complete guide failure where the stars wandered around by 30 arcseconds. Evening twilight flats showed a very large elliptical “Eye of Sauron” pattern which is new; bright in some filters and dark in others. There are also three new narrow features looking almost like scratches. I’ll check in the morning to see how reproducible these flat-field features are and compare sensitivities using Landolt standards with the latest pre-shutdown values. The cooler is struggling even with set point -30 C - the reported value ranged from -22 to -28.
05/27/2020
Xianming Han
none
CCD
7
1.5
Software/Control PC Issues
Observed eclipsing binary. 1. The ACE control cabinet needed rebooting at the beginning of the night. 2. The FWHM was spreading too widely for the same focus position, so that the auto focusing was confused about what is the best focus position. Solution: narrow down the search range. 3. Telescope lost tracking about 2 hours before the end of the night. Once that was fixed, the dome lost tracking It turns out the dome does not move at all, not even manually. So closed the telescope.